Polaski Company manufactures and sells a single product called a Ret. Operating at capacity, the company can produce and sell 36,000 Rets per year. Costs associated with this level of production and sales are given below:

Unit Total
Direct materials $ 15 $ 540,000
Direct labor 8 288,000
Variable manufacturing overhead 3 108,000
Fixed manufacturing overhead 5 180,000
Variable selling expense 4 144,000
Fixed selling expense 6 216,000
Total cost $ 41 $ 1,476,000

The Rets normally sell for $46 each. Fixed manufacturing overhead is $180,000 per year within the range of 27,000 through 36,000 Rets per year.

Required:

1. Assume that due to a recession, Polaski Company expects to sell only 27,000 Rets through regular channels next year. A large retail chain has offered to purchase 9,000 Rets if Polaski is willing to accept a 16% discount off the regular price. There would be no sales commissions on this order; thus, variable selling expenses would be slashed by 75%. However, Polaski Company would have to purchase a special machine to engrave the retail chain’s name on the 9,000 units. This machine would cost $18,000. Polaski Company has no assurance that the retail chain will purchase additional units in the future. What is the financial advantage (disadvantage) of accepting the special order? (Round your intermediate calculations to 2 decimal places.)

2. Refer to the original data. Assume again that Polaski Company expects to sell only 27,000 Rets through regular channels next year. The U.S. Army would like to make a one-time-only purchase of 9,000 Rets. The Army would pay a fixed fee of $1.20 per Ret, and it would reimburse Polaski Company for all costs of production (variable and fixed) associated with the units. Because the army would pick up the Rets with its own trucks, there would be no variable selling expenses associated with this order. What is the financial advantage (disadvantage) of accepting the U.S. Army's special order?

3. Assume the same situation as described in (2) above, except that the company expects to sell 36,000 Rets through regular channels next year. Thus, accepting the U.S. Army’s order would require giving up regular sales of 9,000 Rets. Given this new information, what is the financial advantage (disadvantage) of accepting the U.S. Army's special order?
Answer #1
1 Price per unit for the proposed order $38.64 (46*(1-0.16)
Variable Costs per unit:
Direct materials $15
Direct labor $8
Variable manufacturing overhead $3
Variable Selling Expense=4*(1-0.75) $1
Total Variable Cost per unit $27
Unit Contribution Margin =38.64-27= $11.64
Total Contribution for the order $104,760 (9000*11.64)
Less: Investment in special machine $18,000
Net Profit $86,760
Financial Advantage for accepting the order $86,760
2 US Army Order
Total Cost Per unit $41
Less: Variable Selling expense $4
Cost that will be reimbursed per unit $37
Add: Fixed Fee per unit $1.20
Selling Price per unit $38.20
Variable Cost per unit(27-1) $26 $27 per unit calculted in Problem 1 less $1 Variable selling expense
Unit Contribution Margin =38.20-26 $12.20
Total Contribution Margin=12.2*9000 $109,800
Financial Advantage for accepting the order $109,800
3 IF ORDER IS NOT ACCEPTED
Regular Order:
Selling Price Per unit $46
Variable Cost Per Unit: $30 (15+8+3+4)
Unit Contribution $16
Total Contribution =36000*16= $576,000
IF US ARMY ORDER IS ACCEPTED:
Contribution from Regular Order $432,000 (27000*16)
Contribution from US Army Order $109,800 (12.2*9000)
Total Contribution $541,800
Financial Disadvantage =576000-541800= $34,200